Please solve for me x if log (10^x+2)=2x+1
Found 2 solutions by jorel1380, ikleyn:
Answer by jorel1380(3719)

If log (10^x+2)=2x+1, then:
10^(2x+1)=10^(x+2)
2x+1=x+2
x=1
☺☺☺☺

Answer by ikleyn(52915)

.
It is just the time to learn on how to write correctly parentheses in formulas . . .